there were some birds & deer in a zoo. if heads are counted, they are 200 & if legs are counted they are 580. how many birds are there?
Answer by fractalier(6550)

Call the number of birds and deer, b and d.
From the facts of the problem
b + d = 200
and
2b + 4d = 580
We can solve this a couple of ways, but let's multiply the first equation by two and subtract it from the second...
2b + 4d = 580
-(2b + 2d = 400)
and we get
2d = 180
d = 90 deer and then
b = 110 birds